# Sequential Randomization Tests Using e-values: Applications for trial monitoring

### v5

@@ 7 Discussion
The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a). The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses. It can be extended to continuous endpoints and time-to-event analyses.
@@ 7.1 Operating characteristics
Simulations demonstrate that the method properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
Both the continuous (e-RTC) and the survival (e-survival) implementations had similar characteristics: Power was traded for continuous Type 1 error control. All simulations showed that Type 1 error was exceptionally well controlled but power was nominally low.
These results should be interpreted carefully. From a clinical trial perspective, it is uncertain whether those methods can replace traditional frequentist or Bayesian paradigms, but they may provide a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule. If the e-process crosses its threshold, one could consider stopping the trial, for example. If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
@@ 7.2 What is the null hypothesis being tested?
A useful analogy is a casino. Under the null, the house is fair: no betting strategy can systematically grow wealth. The e-value quantifies accumulated evidence that the game is beatable. Rejecting the null means we have found a profitable strategy—outcomes predict assignments better than chance. The user is also referenced to the pioneer lessons by ramdas2021game.
@@ 7.3 Relationship to existing work
The betting framework for hypothesis testing was developed by shafer2021. E-values and e-processes have been extensively studied (vovk2021; ramdas2022; ramdas2025). pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which applies the betting framework directly to randomized experiments: an analyst sequentially bets on treatment assignments based on observed outcomes, with wealth forming a test martingale under the null. The binary approach implements this framework with a specific adaptive betting strategy tied to outcome values. Betting approaches have been established for estimating means of bounded random variables (waudbysmith2023estimating). The continuous extension adapts these principles to the two-sample randomization setting using a standardization strategy.
koning2025 develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations. pmlr-v146-grunwald21a developed the ‘Safe Log-rank Test’ based on evaluating likelihood ratios with specific priors on the hazard ratio to ensure growth rate optimality. In contrast, the survival approach constructs a linear test martingale directly from the score function using an adaptive betting strategy. Rather than relying on likelihood integration or specific priors, the survival approach process uses a heuristic ‘plug-in’ estimate of the effect direction, modulated by a ramping function. This offers a computationally simple, algorithmic alternative that derives validity strictly from the randomization probabilities within the risk set, without requiring the full apparatus of partial likelihood theory.
@@ 7.4 Limitations
Several limitations should be noted. This is an experimental method under development. Simulations are not exhaustive and were done for binary, continuous endpoints, and time to event. It is uncertain how this method would behave in more complex models, including competing risk models.
The methods proposed test only whether there are differences between arms. They do not directly estimate treatment effects or provide confidence intervals. The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates. Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power. Fourth, our simulations used specific betting strategies; other choices may yield different operating characteristics. Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power. Finally, it is unclear how this method will behave in situations where heterogeneity in treatment effects exist or there are temporal instabilities in effect size. The method could be extended to bet according to relative effect size approaches, such as the odds ratio. This is also under development.
@@ 8.2 LLM use statement
Large language models were extensively used in this work. The author had the idea that perhaps the e-value and e-process machinery could be used to bet against randomization which would result in a continuous trial monitoring tool. They uploaded the references in this manuscript to Gemini 3.0 Pro for brainstorming, which quickly resulted in a preliminary version. This was refined, tested, and debugged using Claude 4.5 Opus and ChatGPT 5.1 Pro. Gemini 3.0 Pro aided with coding for survival approach. Claude 4.5 Opus provided final spellchecking and organization of this draft, which remains suboptimal despite the efforts of several LLMs.
@@ 8.4 Version Control
    Fourth Version (Dec 17, 2025): Correction on signal direction on e-RTC. Text adjustments
@@ Appendix A R Code
compute_eRTC <- function(treatment, outcome, p = 0.5, burn_in = 20,
ramp = 50, c_max = 0.6) {
@@
direction <- if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) 0 else sign(mean_trt - mean_ctrl)
@@
alpha = 0.05, burn_in = 20, ramp = 50,
@@
burn_in = 20, ramp = 50, lambda_max = 0.2) {

### v6

@@ 7.6 Interpretation
8 Betting Strategy Design: The Wage Asymmetry
The five e-RT variants use different betting strategies. Binary and continuous methods use adaptive wagers that learn from data; survival uses a fixed wager $\lambda=0.25$; deaths-only uses adaptive full Kelly. This section explains why these differences are principled, not accidental.
8.1 The Core Asymmetry: Update Density and Over-Betting
The wealth process is a product of multipliers: $W_{n}=\prod_{i=1}^{n}(1+b_{i}u_{i})$. When the bet magnitude $|b_{i}|$ is too large relative to the true effect size (“over-betting”), some multipliers fall well below 1. The critical question is: how fast does wealth recover?
The answer depends on how frequently the product updates:
Survival (e-RTs): Wealth updates only at events (deaths or failures). Between events, wealth is unchanged. A trial with $N=500$ events produces 500 multiplications. Over-betting makes individual multipliers more volatile, but the limited number of updates means wealth bleeds slowly. In simulation, setting $\lambda=0.25$ when the true hazard ratio is 0.90 (where the optimal $\lambda$ is approximately 0.15) yields a median final e-value of approximately 0.04—low, but not zero. The wealth process is damaged but not destroyed.
[…]
Continuous (e-RTc): The problem is even worse. Beyond the dense updates (every patient), the $g$-score introduces a second source of variability in each wager. Fixed-wage betting is catastrophic at all tested levels for effects of $d\leq 0.20$.
In plain language: imagine a gambler who bets too aggressively. If they play once a week (survival), a bad streak hurts but they can recover. If they play every hour (binary), the same over-bet compounds into ruin. If they play every hour and each bet has extra noise on top of the sizing error (continuous), the ruin is faster still.
8.2 Survival: Why Fixed $\lambda=0.25$ Works
For survival data, the bet has fixed magnitude but adaptive direction:
$b_{i}=c_{i}\times\lambda_{\max}\times\text{sign}(Z_{i-1})$ (38)
where $\lambda_{\max}=0.25$ is fixed and $Z_{i-1}$ is the cumulative log-rank score.
Simulation studies across the clinically relevant range of hazard ratios show that $\lambda=0.25$ is not universally optimal but is robustly adequate:
Table 10: Survival: Fixed $\lambda=0.25$ vs. Adaptive Half-Kelly
True HR $|\log(\text{HR})|$ Fixed $\lambda=0.25$ Adaptive $\frac{1}{2}$-Kelly Difference
0.70 0.357 43.1% 26.6% ++16.5pp
0.75 0.288 58.9% 35.0% ++23.9pp
0.80 0.223 62.6% 35.9% ++26.7pp
0.85 0.163 55.2% 37.8% ++17.4pp
0.90 0.105 40.5% 41.6% -1.1pp
Fixed $\lambda=0.25$ outperforms adaptive half-Kelly by 17–27 percentage points across HR =0.70=0.700.850.85. At HR =0.90=0.90 (a small effect), the strategies are approximately equivalent. The adaptive approach converges correctly to the true $\log(\text{HR})$ but produces bets that are systematically too conservative: for HR =0.80=0.80, half-Kelly yields $\lambda\approx 0.11$, which is well below $0.25$.
The fixed strategy succeeds because events are sparse. A trial with approximately 500 events tolerates moderate over-betting without catastrophic wealth destruction. Pre-specifying $\lambda$ is analogous to pre-specifying the alternative hypothesis for sample size calculation—every frequentist trial makes a similar design choice.
8.3 Binary: Why Adaptive Half-Kelly Is Necessary
For binary data, the wager scales with the running risk difference:
$\lambda_{i}=0.5\pm 0.5\times c_{i}\times\hat{\delta}_{i-1}$ (39)
The factor of $0.5$ before $c_{i}\hat{\delta}$ implements half-Kelly betting. Why not use the full estimate, or a fixed wager?
Simulation demonstrates the catastrophic cost of over-betting in binary trials. For a true ARR of 5% (Kelly-optimal $\lambda\approx 0.05$):
Table 11: Binary: Cost of Over-Betting (True ARR = 5%, N =2,942=2{,}942)
Fixed $\lambda$ Relative to Kelly Power Median Final E-value
0.05 $1\times$ 56.3% 1.15
0.10 $2\times$ 21.9% $\approx 0$
0.15 $3\times$ 16.5% $\approx 0$
0.20 $4\times$ 10.0% $\approx 0$
At $2\times$ Kelly, the median e-value is already zero—wealth is irreversibly destroyed. Every patient updates the product, and thousands of slightly-wrong multiplications compound into oblivion.
Moreover, no single fixed $\lambda$ works universally. The optimal fixed $\lambda$ ranges from 0.05 for small effects to 0.15 for moderate effects. In practice, trialists routinely power for optimistic effects (e.g., ARR $10\%$) when the truth may be 3–5%. If $\lambda$ inherits this optimism, wealth is destroyed.
The adaptive half-Kelly wager is not a power sacrifice—it is insurance against the trialist’s optimism. The 7-percentage-point power loss relative to a perfectly calibrated fixed wager ($1\times$ Kelly: 56.3% vs. adaptive: 49.4%) is the cost of being agnostic to the design alternative. That cost is worth paying because the alternative (inheriting an inflated design effect) risks irreversible wealth destruction.
8.4 Continuous: The Strongest Case for Adaptive Betting
The continuous e-RTc has a “doubly adaptive” structure:
$\lambda_{i}=0.5+c_{i}\times c_{\max}\times g_{i}\times\hat{d}_{\text{past}}$ (40)
Both $g_{i}$ (observation-level informativeness) and $\hat{d}_{\text{past}}$ (running Cohen’s $d$) are adaptive. Removing the magnitude of $\hat{d}$ and using only its sign (making $\lambda$ “fixed” in the direction sense) is catastrophic:
Table 12: Continuous: Adaptive vs. Fixed Direction Magnitude
Cohen’s $d$ NN Adaptive Fixed $c=0.3$ Fixed $c=0.6$
0.50 128 16.8% 15.5% 45.3%
0.30 352 50.1% 50.1% 25.2%
0.20 788 53.9% 29.3% 13.6%
At Cohen’s $d=0.20$ (a typical clinical trial effect size), fixed $c=0.6$ retains only 13.6% power (vs. 53.9% adaptive), and fixed $c=0.3$ drops to 29.3%. The median final e-value is $\approx 0$ in both cases—wealth is irreversibly destroyed. The $g$-score adds a second noise source that, combined with dense updates, makes even moderate fixed betting lethal.
8.5 The Design Principle
Table 13: Summary: Betting Strategy by e-RT Variant
Property e-RTs e-RTd e-RT e-RTc
Update frequency Events only Events only Every patient Every patient
Over-bet cost Low (slow bleed) Moderate Catastrophic Catastrophic
Universal $\lambda$ possible? Yes Marginal No No
Kelly fraction Fixed ($\lambda=0.25$) Full Half Doubly adaptive
The hierarchy of over-betting risk is: continuous >> binary $\gg$ survival. This maps directly to how frequently the wealth product is updated. More frequent updates mean more opportunities for over-betting to compound, requiring increasingly conservative or adaptive strategies.
The e-RTd falls between survival and binary. Like survival, it updates only at events (deaths), making the product length moderate. This permits full Kelly rather than half-Kelly. However, unlike survival where $\lambda_{\max}=0.25$ is fixed, e-RTd uses the plug-in $\hat{p}$ directly because the death-coin estimator converges quickly and the natural scale of the problem ($p\in[0,1]$) bounds the wager without external calibration.
@@ 9 Discussion
The e-RT family comprises five nonparametric sequential tests for randomized trials based on the betting framework for e-values (i-bet Duan et al. (2022)). All variants require only that treatment assignment is randomized—no distributional assumptions about outcomes are needed. This makes them robust complements to model-based analyses. The variants cover binary outcomes, deaths-only monitoring, continuous endpoints, time-to-event analyses, and multi-state trajectory data.
@@ 9.1 Operating characteristics
Across all five variants, simulations demonstrate proper Type I error control (consistently 2–4% vs. nominal 5%), confirming the theoretical guarantee from the martingale property. Power varies by variant and scenario: for binary outcomes, approximately 50% power for early stopping in trials designed with 80% frequentist power, and 63–66% in trials designed with 90% power. Continuous and survival variants show similar patterns. The multi-state variant achieved 89% power at N=1, [^0] N=1{,}000 for the specific transition matrix tested. The deaths-only variant (e-RTd) requires approximately $2.5\times$ sample size inflation compared with the frequentist benchmark, the cost of discarding survivor information.
When early stopping occurs across all variants, it typically happens at approximately 45–56% of the planned sample size or event count.
These results should be interpreted carefully. From a clinical trial perspective, it is uncertain whether these methods can replace traditional frequentist or Bayesian paradigms, but they may provide a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule. If the e-process crosses its threshold, one could consider stopping the trial. If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
9.2 Traditional statistics at crossing
When an e-RT crosses its threshold ($\geq 1/\alpha$), one may ask what a traditional frequentist analysis would report at that exact moment. In practice, the standard test will agree: the p-value will be small and the confidence interval will exclude the null. The e-RT is not detecting phantoms.
However, there is a critical distinction. The traditional analysis at crossing is formally invalid because the analyst observed the data at multiple time points and stopped at a favorable one. The frequentist p-value is anti-conservative (too small) and the standard confidence interval has less than 95% coverage at this data-driven stopping time. The e-value, by contrast, is valid regardless of the stopping rule—this is the entire point of the anytime-valid guarantee. The e-value itself carries the inferential weight; the traditional statistics at crossing serve only as a sanity check confirming that the e-RT is detecting a real signal, not an artifact.
@@ 9.3 What is the null hypothesis being tested?
A useful analogy is a casino. Under the null, the house is fair: no betting strategy can systematically grow wealth. The e-value quantifies accumulated evidence that the game is beatable. Rejecting the null means we have found a profitable strategy—outcomes predict assignments better than chance. The user is also referenced to the pioneer lessons by Ramdas (2021).
@@ 9.4 Relationship to existing work
The betting framework for hypothesis testing was developed by Shafer (2021). E-values and e-processes have been extensively studied (Vovk and Wang, 2021; Ramdas et al., 2022; Ramdas and Wang, 2025). Duan et al. (2022) introduced interactive rank testing by betting (i-bet), which applies the betting framework directly to randomized experiments: an analyst sequentially bets on treatment assignments based on observed outcomes, with wealth forming a test martingale under the null. The binary approach implements this framework with a specific adaptive betting strategy tied to outcome values. Betting approaches have been established for estimating means of bounded random variables (Waudby-Smith and Ramdas, 2023). The continuous extension adapts these principles to the two-sample randomization setting using a standardization strategy.
Koning (2025) develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations. Grünwald et al. (2021) developed the ‘Safe Log-rank Test’ based on evaluating likelihood ratios with specific priors on the hazard ratio to ensure growth rate optimality. In contrast, the survival approach constructs a linear test martingale directly from the score function using an adaptive betting strategy. Rather than relying on likelihood integration or specific priors, the survival approach process uses a heuristic ‘plug-in’ estimate of the effect direction, modulated by a ramping function. This offers a computationally simple, algorithmic alternative that derives validity strictly from the randomization probabilities within the risk set, without requiring the full apparatus of partial likelihood theory.
@@ 9.5 Limitations
Several limitations should be noted. This is an experimental method under development. Simulations are not exhaustive, and the operating characteristics reported are specific to the scenarios tested. It is uncertain how these methods would behave in more complex models, including competing risk models.
The methods test only whether there are differences between arms; they do not provide point estimates or confidence intervals. The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates. For trials where parametric assumptions are plausible, model-based sequential methods will generally have better power. Our simulations used specific betting strategies; other choices may yield different operating characteristics. The betting strategy design section provides guidance on strategy selection, but optimal calibration for specific clinical scenarios remains an open question. Finally, it is unclear how these methods will behave when heterogeneity in treatment effects exists or there are temporal instabilities in effect size. Extensions to relative effect size approaches (e.g., odds ratio) are under development.
@@ 10.2 LLM use statement
Large language models were extensively used in this work. The author had the idea that perhaps the e-value and e-process machinery could be used to bet against randomization which would result in a continuous trial monitoring tool. They uploaded the references in this manuscript to Gemini 3.0 Pro for brainstorming, which quickly resulted in a preliminary version. This was refined, tested, and debugged using Claude 4.5 Opus and ChatGPT 5.1 Pro. Gemini 3.0 Pro aided with coding for survival approach. Claude Opus 4.6 aided with the deaths-only extension and the wage asymmetry analysis in V6.
@@ 10.4 Version Control
    Fourth Version (Dec 17, 2025): Correction on signal direction on e-RTC. Text adjustments.
@@
  6. 6.
    Sixth Version (Feb 15, 2026): Added deaths-only monitoring (e-RTd). Added betting strategy design section explaining wage asymmetry across variants. Reconciled e-RTc parameters (burn-in 50, ramp 100) and direction formula (full Cohen’s $d$ clamped to [1,1][-1,1]). Added traditional statistics at crossing discussion. Updated abstract and introduction to cover all five variants. Reordered sections.
@@ Appendix A R Code
# --- e-RTd: Deaths-Only Monitoring ---
compute_eRTd <- function(arms, burn_in = 30, ramp = 50, threshold = 20) {
n <- length(arms)
if (is.factor(arms)) arms <- as.numeric(arms) - 1
arms <- as.numeric(arms)
wealth <- numeric(n)
wealth[1] <- 1
d_trt <- 0; d_ctrl <- 0
crossed <- FALSE; crossed_at <- NA
for (i in 1:n) {
total <- d_trt + d_ctrl
p_obs <- if (total > 0) d_trt / total else 0.5
if (i > burn_in && total > 0) {
c_i <- min(1, max(0, (i - burn_in) / ramp))
lambda <- 0.5 + c_i * (p_obs - 0.5)
lambda <- max(0.001, min(0.999, lambda))
} else {
lambda <- 0.5
mult <- if (arms[i] == 1) lambda / 0.5 else (1 - lambda) / 0.5
wealth[i] <- if (i == 1) mult else wealth[i - 1] * mult
if (arms[i] == 1) d_trt <- d_trt + 1 else d_ctrl <- d_ctrl + 1
if (!crossed && wealth[i] >= threshold) {
crossed <- TRUE; crossed_at <- i
total <- d_trt + d_ctrl
final_p <- if (total > 0) d_trt / total else 0.5
final_rr <- if (final_p > 0.001 && final_p < 0.999) {
final_p / (1 - final_p)
} else if (final_p <= 0.001) { 0 } else { Inf }
list(wealth = wealth, crossed = crossed, crossed_at = crossed_at,
final_p = final_p, final_rr = final_rr, d_trt = d_trt, d_ctrl = d_ctrl)
simulate_deaths <- function(n_deaths, p_trt_given_death) {
rbinom(n_deaths, 1, p_trt_given_death)
death_coin <- function(p_ctrl, p_trt) {
p_trt / (p_trt + p_ctrl)
expected_deaths <- function(n_patients, p_ctrl, p_trt) {
ceiling(n_patients / 2 * (p_ctrl + p_trt))
simulate_eRTd <- function(n_sims = 1000, p_ctrl, p_trt,
n_patients = NULL, target_power = 0.80,
alpha = 0.05, burn_in = 30, ramp = 50) {
threshold <- 1 / alpha
if (is.null(n_patients)) {
$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_trt,
power = target_power, sig.level = alpha)
n_freq <- 2 * ceiling(ss\$n)
n_patients <- ceiling(n_freq * 2.5)
n_deaths_null <- expected_deaths(n_patients, p_ctrl, p_ctrl)
n_deaths_alt <- expected_deaths(n_patients, p_ctrl, p_trt)
p_coin_alt <- death_coin(p_ctrl, p_trt)
null_rej <- 0; alt_rej <- 0
alt_crossing <- numeric(n_sims)
for (sim in 1:n_sims) {
# Null
deaths <- simulate_deaths(n_deaths_null, 0.5)
res <- compute_eRTd(deaths, burn_in, ramp, threshold)
if (res\$crossed) null_rej <- null_rej + 1
# Alternative
deaths <- simulate_deaths(n_deaths_alt, p_coin_alt)
res <- compute_eRTd(deaths, burn_in, ramp, threshold)
if (res\$crossed) alt_rej <- alt_rej + 1
alt_crossing[sim] <- ifelse(res\$crossed, res\$crossed_at, NA)
list(type1 = null_rej / n_sims, power = alt_rej / n_sims,
median_crossing = median(alt_crossing, na.rm = TRUE),
n_patients = n_patients, n_deaths_alt = n_deaths_alt)
plot_trajectories_eRTd <- function(n_trials = 30, n_deaths,
p_trt_given_death,
burn_in = 30, ramp = 50, alpha = 0.05,
title = NULL) {
threshold <- 1 / alpha
if (is.null(title)) {
title <- if (p_trt_given_death == 0.5) {
sprintf("e-RTdUnderNull(p=0.50,%ddeaths)", n_deaths)
} else {
sprintf("e-RTdUnderAlternative(p=%.3f,%ddeaths)",
p_trt_given_death, n_deaths)
trajectories <- lapply(1:n_trials, function(i) {
deaths <- simulate_deaths(n_deaths, p_trt_given_death)
res <- compute_eRTd(deaths, burn_in, ramp, threshold)
data.frame(death = 1:n_deaths, wealth = res\$wealth, trial = i)
df <- bind_rows(trajectories)
ggplot(df, aes(x = death, y = wealth, group = trial)) +
geom_line(alpha = 0.4, color = "steelblue") +
geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
scale_y_log10() +
labs(title = title, x = "DeathNumber", y = "e-value(logscale)") +
annotate("text", x = n_deaths * 0.95, y = threshold * 1.5,
label = sprintf("1/alpha=%.0f", threshold),
color = "red", hjust = 1) +
theme_minimal() +
theme(plot.title = element_text(face = "bold"),
panel.grid.minor = element_blank())
signal_concentration_table <- function(
baseline_rates = c(0.10, 0.15, 0.20, 0.25, 0.30, 0.35, 0.40),
arr = 0.05) {
data.frame(
baseline = baseline_rates,
trt_rate = baseline_rates - arr,
death_coin = sapply(baseline_rates, function(p) death_coin(p, p - arr)),
tilt_from_half = sapply(baseline_rates,
function(p) abs(death_coin(p, p - arr) - 0.5)),
arr = arr
@@
compute_eRTC <- function(treatment, outcome, p = 0.5, burn_in = 50,
ramp = 100, c_max = 0.6) {
@@
# Cohen’s d (clamped to [-1, 1]) -- not just sign
if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) {
d_hat <- 0
} else {
sd_trt <- sd(out_prev[trt_prev == 1])
sd_ctrl <- sd(out_prev[trt_prev == 0])
if (is.na(sd_trt) $|$ sd_trt == 0) sd_trt <- 1
if (is.na(sd_ctrl) $|$ sd_ctrl == 0) sd_ctrl <- 1
s_pooled <- sqrt((sd_trt^2 + sd_ctrl^2) / 2)
d_hat <- max(-1, min(1, (mean_trt - mean_ctrl) / s_pooled))
@@
alpha = 0.05, burn_in = 50, ramp = 100,
@@
burn_in = 30, ramp = 50, lambda_max = 0.25) {